It takes Jack 15 hours longer to drain a basement than Jill. If they work together, it takes 4 hours to drain a basement. How long does it take for each person to working alone?
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

It takes Jack 15 hours longer to drain a basement than Jill. If they work together, it takes 4 hours to drain a basement. How long does it take for each person to working alone?
--------------
Draining the basement is the job.
Jill does it in some number, J, of hours, so she does 1/J per hour.
Jack does it in J+15 hours, so he does 1/(J+15) per hour.
Together, they do 1/J + 1/(J+15) per hour, and it takes them 4 hours.
Together, the do 1/4 of the job.
---> 1/J + 1/(J+15) = 1/4
Can you solve for J ?

Answer by stanbon(75887)

It takes Jack 15 hours longer to drain a basement than Jill. If they work together, it takes 4 hours to drain a basement. How long does it take for each person to working alone?
Jack DATA:
time = x + 15 hrs/job ; rate = 1/(x+15) job/hr
----
Jill DATA:
time = x hrs/job ; rate = 1/x job/hr
------------------------
Together DATA:
time = 4 hrs/job ; rate = 1/4 job/hr
------------------------------------------
Equation:
rate + rate = together rate
1/x + 1/(x+15) = 1/4
------------------------------
Multiply thru by 4x(x+15) to get::
4(x+15) + 4x = x(x+15)
4x + 60 + 4x = x^2 + 15x
----------
x^2 + 7x - 60 = 0
------
(x+12)(x-5) = 0
-----
Positive solution:
x = 5 hrs (Jill's time)
x+15 = 20 hrs (Jack's time)
------------------
Cheers,
Stan H.
------------------